# Genie Sim 3.0 : A High-Fidelity Comprehensive Simulation Platform for Humanoid Robot

### v1

@@ I INTRODUCTION
    We open-source Genie Sim 3.0 simulation platform which integrates scene generation, evaluation synthesis, data acquisition, and closed-loop model assessment. To foster community progress, we open-source 5140 assets, 10,000+ hours dataset, 100,000+ evaluation scenarios, and the complete codebase. By providing massive-scale synthetic dataset and evaluation scenes for insightful capability diagnostics and panoramic evaluation to accelerate model iteration, Genie Sim 3.0 reduces reliance on physical hardware, enhances research and development efficiency, and propels the development of practical embodied AI applications.
@@ III-E Closed-loop Evaluation
  • Integration with common VLA models ($\pi_{0.5}$ [undefau], GO-1 [undefs], UniVLA [undefav], RDT [undefaw], X-VLA [undefax], etc.)
@@ V Experiment
TABLE I: Experiment Results: four tasks, each with four models tested in both real and sim environments
@@
1500 eps sim 0.86 0.85 0.93 0.94 0.80 0.71 0.52 0.60
The validity and reliability of simulation benchmarks and datasets constitute a foundational consideration in embodied intelligence research.
Figure 6: Sim to Real Experiments. Comparison of initial task setup between real and sim testing environments
Two central questions are examined: first, whether a simulation benchmark can evaluate model performance as reliably as physical environment; and second, to what extent data synthetic data can effectively substitute for physically-collected data in model training. To address these questions systematically, a series of experiments are designed to assess the validity and efficacy of Genie Sim 3.0 benchmark and dataset.
V-A Experimental Setup
In this work, we conducted 32 groups of experiments to compare the task success rates of models fine-tuned with varying amounts of real-world and synthetic data, tested in both physics and simulation environments. The $\pi_{0.5}$ model [undefau]— a widely recognized baseline with demonstrated performance in related research— is selected as the base model and Agibot G1 robot is deployed for task execution.
Since the primary objective of these experiments is to evaluate the validity and reliability of the simulation benchmark and dataset, tasks with either excessively high or low success rates are avoided to better reveal performance differences across dataset and environments. In accordance with this principle and to ensure assessment covering both atomic skills and cognitive capabilities, four representative tasks are selected for evaluation:
  1. 1.
    Select Color,
  2. 2.
    Recognize Size,
  3. 3.
    Grasp Targets,
  4. 4.
    Organize Items.
For each task, models are post-trained with four different setups:
  1. 1.
    200 episodes of real-env data,
  2. 2.
    500 episodes of real-env data,
  3. 3.
    500 episodes of sim-env data,
  4. 4.
    1500 episodes of sim-env data.
Each experiment is evaluated over 50 trials in real world environment and over 250 trails in simulation.
Figure 7: Model Performance on Real-World Environment. The success rates of the 4 tasks, each with four training setups are shown in the figure with different colors.
V-B Results and Analysis
Figure 8: Performance Comparison between Real and Sim Environments. The performance in simulation share the same trend with it in real-world environment.
Effectiveness of the Synthetic Data. In general, the success rates of the tasks significantly improve with an increasing volume of training data—for both real and synthetic data—which aligns well with the scaling laws in learning performance, see Fig. 7. Increased task complexity leads to reduced success rates for models trained on datasets of identical size. Most importantly, the model trained exclusively on 1500 eps of synthetic data attains the highest zero‑shot success rates across all four tasks in the real‑world environment, as shown in Tab. I. When comparing data of the same quantity (500 eps), models trained on real‑robot data consistently outperform those trained on synthetic data. This advantage can be attributed to the higher physical fidelity of real data, which inherently captures complex real‑world phenomena such as friction, contact dynamics, and collisions—aspects that are often imperfectly modeled and lead to sim‑to‑real gaps in synthetic datasets.
Figure 9: Correlation Analysis of Model Performance on Sim and Real Environments. All 16 models are evaluated both in sim and real environments, which are tagged on the right side of the figure, each with a different color and marker.
Figure 10: Success Rate on both Sim and Real Env in Heatmap. The comparison of the left and right charts intuitively reflects the consistency between the real and the sim environment.
Nevertheless, real‑world robotics data collection incurs substantial costs and time constraints, limiting the scale and diversity of available datasets. In contrast, synthetic data can be generated efficiently at scale, allowing for extensive randomization in parameters such as object textures, lighting conditions, physics properties, and task variations. By scaling the synthetic dataset to 1500 episodes and systematically varying simulation conditions, the resulting training distribution effectively mitigates the domain gap, thereby enhancing the model’s generalization capabilities to real‑world settings.
Thus, the large‑scale, systematically randomized synthetic data pipeline implemented in GenieSim 3.0 offers a pragmatic and scalable alternative to real‑robot data collection, enabling the training of robust models capable of high zero‑shot performance in physical environments.
Consistency between Sim and Real Environments. Fig. 8 provides a comparative analysis of model performance in simulated versus physical environments. While minor discrepancies are observed across tasks—primarily attributable to systematic differences from approximations of contact dynamics and friction in simulation, as well as stochastic factors inherent to real-world deployment such as unstructured environmental perturbations—the overall performance trends remain consistent. This close correspondence is further illustrated in the success-rate heatmaps presented in the Fig. 10. A quantitative correlation analysis, summarized in the scatter plot of the Fig. 9, reveals a statistically significant strong positive relationship between simulated and real-world performance ($R^{2}=0.924$), with a best-fit slope $\approx 1.045$. These results indicate that evaluations conducted in Genie Sim 3.0 yield reliable conclusions that are consistent with trends observed in real-world testing.
Given this high level of alignment, alongside the advantages of simulation—including fully automated evaluation pipelines, high reproducibility, scalable scenario generation, support for concurrent batch executions, and significantly lower operational costs—Genie Sim 3.0 serves as a robust and efficient platform for model assessment. Its use can substantially accelerate research iteration and facilitate experimental scaling beyond the practical constraints of physical testing.
@@ ACKNOWLEDGMENT
We would like to express our sincere gratitude to Sukai Wang and Xiaowei Cai for their assistance with model inference and training; to Haowen Yin, Yanping Zhou, Zengcheng Zhou, and Cheng Ruan for their support in experimental testing; to Chengyue Zhao, Xinke Yu, and Jialu Li for their work on publicity and design; and to Cheng Jing, Haoyu Cao, Chi Zhang, Guangte Xiang, Jun Dai for their help with data processing.

### v2

@@ I INTRODUCTION
    We open-source Genie Sim 3.0 simulation platform which integrates scene generation, evaluation synthesis, data acquisition, and closed-loop model assessment. To foster community progress, we open-source 5,140 assets, 10,000+ hours dataset, 100,000+ evaluation scenarios, and the complete codebase. By providing massive-scale synthetic dataset and evaluation scenes for insightful capability diagnostics and panoramic evaluation to accelerate model iteration, Genie Sim 3.0 reduces reliance on physical hardware, enhances research and development efficiency, and propels the development of practical embodied AI applications.
@@ III-E Closed-loop Evaluation
  • Integration with common VLA models ($\pi_{0.5}$ [undefau], GO-1 [undefs], GR00T, UniVLA [undefav], RDT [undefaw], X-VLA [undefax], etc.)
@@ V Experiment
To systematically assess the validity and utility of Genie Sim as a simulation benchmark and data generation platform, we introduce four closed-loop evaluation suites. GenieSim-Sim2Real is designed to validate the data quality of Genie Sim and the fidelity of its simulated environments: it examines whether large-scale synthetic data can match or exceed real-world data in model training, and whether benchmark evaluations in simulation yield conclusions consistent with physical deployment. GenieSim-Instruction focuses on language-conditioned manipulation, probing how well policies can follow diverse and compositional natural language instructions. GenieSim-Robust systematically evaluates policy robustness by introducing controlled perturbations across robot configuration, camera parameters, control latency, and environmental conditions. GenieSim-Manipulation covers a broad set of challenging manipulation tasks that require precise motor skills, long-horizon planning, and dexterous object interactions. Together, these four suites provide complementary perspectives on policy capabilities and benchmark reliability.
Figure 6: GenieSim-Sim2Real Task Setups. Comparison of initial task configurations between real and simulated testing environments across all evaluated tasks. Rows 1 and 3 show real-world environments, while rows 2 and 4 show their corresponding simulated counterparts.
V-A Experimental Setup
GenieSim-Sim2Real. All experiments within this suite use the $\pi_{0.5}$ model [undefau] as the base policy, post-trained on varying amounts of real-world or synthetic data. The Agibot G1 robot is deployed for physical task execution. Each training configuration is evaluated over 50 trials in real-world environments and in simulation, ensuring direct comparability between the two evaluation modalities.
GenieSim-Instruction, GenieSim-Robust, and GenieSim-Manipulation. For these three benchmark suites, three state-of-the-art vision-language-action models are compared: $\pi_{0.5}$ [undefau], GR00T-N1.6 [undefay], and $\pi_{0}$ [undefaz]. Each model is fine-tuned with task-specific synthetic data generated by GenieSim and evaluated on the Agibot G2 robot using GenieSim’s automated closed-loop evaluation pipeline.
V-B GenieSim-Sim2Real
V-B1 Synthetic Data Scaling
TABLE I: Synthetic Data Scaling Results. Task success rates of $\pi_{0.5}$ fine-tuned with varying amounts of real and synthetic data, evaluated in both simulated and real environments.
@@ V-B1 Synthetic Data Scaling
1500 eps sim 0.86 0.85 0.93 0.94 0.72 0.71 0.52 0.60
We conduct 32 groups of experiments across four representative tasks: Select Color, Recognize Size, Grasp Targets, and Organize Items. These tasks span a range of complexities—from atomic manipulation skills to language-conditioned cognitive reasoning—while avoiding floor or ceiling effects that would obscure performance differences. For each task, $\pi_{0.5}$ is fine-tuned under four data configurations: 200 and 500 episodes of real-world data, and 500 and 1500 episodes of synthetic data. Results are reported in Tab. I.
Effectiveness of Synthetic Data. Success rates improve monotonically with increasing data volume for both real and synthetic sources, consistent with scaling laws in robot learning. At equivalent scale (500 eps), models trained on real-world data outperform those trained on synthetic data, reflecting the higher physical fidelity of real demonstrations. However, scaling synthetic data to 1500 episodes enables the model to surpass all real-data baselines across all four tasks in zero-shot real-world evaluation. This demonstrates that Genie Sim’s systematic domain randomization—spanning object textures, lighting conditions, physics parameters, and task variations—effectively bridges the domain gap at scale, offering a practical and scalable alternative to real-world data collection.
Consistency between Sim and Real Environments. Model performance across simulated and physical environments is compared. Despite minor discrepancies attributable to imperfect simulation of contact dynamics and stochastic real-world perturbations, overall performance trends remain consistent across environments. A quantitative correlation analysis (Fig. 7) yields $R^{2}=0.94$ with a slope $\approx 1.025$, indicating that evaluations of Genie Sim reliably predict real-world performance trends and can serve as a cost-efficient substitute for physical benchmarking.
Figure 7: Correlation Analysis of Sim and Real Performance. All 16 model configurations show a strong linear relationship between simulated and real-world success rates ($R^{2}=0.94$).
V-B2 Sim-to-Real Transfer
[…]
Tasks Sim Env Real Env
Sim Data Real Data Sim Data Real Data
(sim-to-sim) (real-to-sim) (sim-to-real) (real-to-real)
Select Color 0.86 0.75 0.85 0.73
Recognize Size 0.93 0.75 0.94 0.75
Grasp Targets 0.72 0.54 0.71 0.58
Organize Items 0.52 0.45 0.60 0.40
Pack in Supermarket 0.94 1.00 0.95 0.95
Sort Fruit 0.90 0.90 1.00 1.00
Place Block into Drawer 0.80 0.90 0.85 0.90
Bimanual Chip Handover 0.80 0.70 0.73 0.71
Avg. 0.81 0.75 0.83 0.75
$\dagger$ Sim Data: $500\sim 1500$ episodes of simulation data. Real Data: $500$ episodes of real-world data.
To further validate the cross-domain transfer capability of Genie Sim, we extend the evaluation to 8 tasks in Tab. II, adding four newly introduced tasks—Pack in Supermarket, Sort Fruit, Place Block into Drawer, and Bimanual Chip Handover—to the original four. For each task, models trained on synthetic data (50050015001500 episodes) are compared against models trained on 500 episodes of real-world data, with both evaluated in simulated and physical environments.
Sim-to-Real. Across all 8 tasks, models trained solely on synthetic data achieve competitive or superior real-world performance relative to models trained on real data, with an average sim-to-real success rate of 0.83 versus 0.75 for real-to-real. This confirms that synthetic data generated by Genie Sim is of sufficient fidelity for direct real-world deployment.
Real-to-Sim. Conversely, models trained on real-world data also transfer well to simulated evaluation, achieving an average of 0.75. This bidirectional transferability demonstrates that the Genie Sim environment faithfully captures the physical structures and task dynamics of real-world settings, establishing it as a reliable evaluation platform.
Taken together, the 32-group scaling experiments and the 8-task transfer study demonstrate that Genie Sim provides both high-quality synthetic data and a reliable simulation benchmark—the two foundational properties required for the credible cross-model evaluations conducted in the following suites.
V-C Genie Sim Benchmark Evaluation
Grounded in the validated reliability of Genie Sim, we present a systematic comparison of $\pi_{0.5}$, GR00T-N1.6, and $\pi_{0}$ across three complementary benchmark suites. Each suite targets a distinct capability dimension, and the high sim-to-real consistency established in Sec. V-B ensures that simulation-based rankings are predictive of real-world policy performance.
GenieSim-Instruction. This suite comprises 10 tasks that require policies to interpret and execute diverse natural language instructions, including directives based on object number, shape, type, color, size, and commonsense or logical reasoning. The tasks are designed to stress-test the language grounding capabilities of VLA models beyond simple keyword matching. As shown in Tab. III, $\pi_{0.5}$ achieves an average success rate of 0.67, substantially outperforming GR00T-N1.6 (0.40) and $\pi_{0}$ (0.28). The performance gap widens on semantically complex tasks such as pick_common_sense (0.35 vs. 0.12 vs. 0.05), indicating that $\pi_{0.5}$’s stronger language model backbone provides a meaningful advantage in instruction-following scenarios.
TABLE III: Experimental results on GenieSim-Instruction. Success rates across 10 language-conditioned manipulation tasks.
Tasks $\boldsymbol{\pi_{0.5}}$ GR00T-N1.6 $\boldsymbol{\pi_{0}}$
pick_block_number 0.73 0.28 0.17
pick_block_shape 0.41 0.15 0.17
pick_common_sense 0.35 0.12 0.05
pick_follow_logic_or 0.58 0.56 0.26
pick_object_type 0.81 0.56 0.27
pick_specific_object 0.58 0.35 0.16
straighten_object 0.66 0.33 0.46
pick_billiards_color 0.81 0.37 0.47
pick_block_color 0.88 0.71 0.40
pick_block_size 0.89 0.52 0.36
Avg. 0.67 0.40 0.28
TABLE IV: Experimental results on GenieSim-Robust. Success rates under 12 categories of environmental and control perturbations, with $\Delta$ denoting the change relative to the Reference condition. The significant drops per model are shown in bold.
Generalization $\boldsymbol{\pi_{0.5}}$ GR00T-N1.6 $\boldsymbol{\pi_{0}}$
Score $\Delta$ Score $\Delta$ Score $\Delta$
Reference 0.92 0.58 0.46
Instruction 0.89 -0.03 0.47 -0.11 0.32 -0.14
Robot Init Base 0.83 -0.09 0.59 +0.01 0.32 -0.14
Robot Init Joint 0.70 -0.22 0.39 -0.19 0.34 -0.12
Robot End Effector 0.42 -0.50 0.30 -0.28 0.26 -0.20
Control Delay 0.76 -0.16 0.57 -0.01 0.40 -0.06
Camera Frame Drop 0.83 -0.09 0.28 -0.30 0.19 -0.27
Camera Noise 0.89 -0.03 0.59 -0.01 0.34 -0.12
Camera Occlusion 0.93 +0.01 0.59 +0.01 0.41 -0.05
Camera Extrinsic 0.39 -0.53 0.27 -0.31 0.22 -0.24
Ambient Lighting 0.85 -0.07 0.54 -0.04 0.44 -0.02
Background 0.90 -0.02 0.57 -0.01 0.40 -0.06
Avg. 0.77 0.48 0.34
TABLE V: Experimental results on GenieSim-Manipulation. Success rates across 10 complex manipulation tasks.
Tasks $\pi_{0.5}$ GR00T-N1.6 $\pi_{0}$
Open Door 0.60 0.35 0.46
Hold Pot 0.35 0.00 0.14
Pour Workpiece 0.95 0.95 0.72
Stock and Straighten Shelf 0.37 0.15 0.21
Take Wrong Item Shelf 0.95 0.65 0.80
Scoop Popcorn 0.78 0.80 0.68
Clean the Desktop 0.16 0.01 0.08
Place Block into Box 0.50 0.30 0.38
Sorting Packages 0.45 0.14 0.13
Sorting Packages Continuous 0.16 0.03 0.00
Avg. 0.53 0.34 0.36
GenieSim-Robust. This suite systematically evaluates policy robustness by introducing 12 categories of controlled perturbations to a reference pick-and-place task, spanning robot initial pose, end-effector configuration, control delay, camera frame drop, image noise, occlusion, extrinsic calibration error, ambient lighting, and scene background. These conditions model realistic deployment variabilities that are difficult to enumerate exhaustively in physical experiments but are readily programmable in simulation.
As reported in Tab. IV, $\pi_{0.5}$ achieves the highest average success rate of 0.77 across all perturbation categories, compared to 0.48 for GR00T-N1.6 and 0.34 for $\pi_{0}$. The relative degradation from the Reference condition is smallest for $\pi_{0.5}$, indicating that it not only achieves higher absolute performance but also degrades more gracefully under distribution shift.
Looking at specific perturbation categories, $\pi_{0.5}$ exhibits the largest drops under Camera Extrinsic ($\Delta=-0.53$) and Robot End Effector ($\Delta=-0.50$) conditions, suggesting sensitivity to physical calibration errors and changes in end-effector configuration. Robot Init Joint also causes a notable decline ($\Delta=-0.22$). In contrast, $\pi_{0.5}$ shows strong robustness to perceptual perturbations such as Camera Occlusion ($\Delta=+0.01$), Camera Noise ($\Delta=-0.03$), and scene Background changes ($\Delta=-0.02$), indicating effective visual invariance to appearance-level disturbances.
For GR00T-N1.6, the most significant performance drops occur under Camera Extrinsic ($\Delta=-0.31$), Camera Frame Drop ($\Delta=-0.30$), and Robot Init Joint ($\Delta=-0.19$), revealing particular vulnerability to temporal input degradation and geometric miscalibration. Meanwhile, GR00T-N1.6 maintains relatively stable performance under Control Delay ($\Delta=-0.01$), Camera Noise ($\Delta=-0.01$), and Background ($\Delta=-0.01$), demonstrating reasonable tolerance to low-level signal perturbations.
$\pi_{0}$ suffers the greatest degradation from Camera Frame Drop ($\Delta=-0.27$) and Camera Extrinsic ($\Delta=-0.24$), followed by Robot End Effector ($\Delta=-0.20$), indicating that its visual perception and spatial reasoning are most susceptible to input incompleteness and calibration shifts. On the other hand, $\pi_{0}$ shows comparatively better resilience to Ambient Lighting ($\Delta=-0.02$) and Camera Occlusion ($\Delta=-0.05$), suggesting some degree of robustness to illumination and partial visual obstruction.
Across all three models, Camera Extrinsic and Robot End Effector perturbations consistently induce the largest performance drops, pointing to these as the most challenging axes of distributional shift for current visuomotor policies.
GenieSim-Manipulation. This suite consists of 10 tasks that target diverse and challenging manipulation skills, including articulated object interaction (Open Door), bimanual collaboration (Hold Pot), granular material handling (Scoop Popcorn), precise placement and sorting (Place Block into Box, Sorting Packages), desktop cleaning, and multi-step shelf organization. These tasks collectively probe the breadth and depth of a policy’s physical manipulation capabilities in realistic scenarios. As shown in Tab. V, $\pi_{0.5}$ achieves the highest average success rate of 0.53, ahead of $\pi_{0}$ (0.36) and GR00T-N1.6 (0.34). High-performance tasks such as Pour Workpiece and Take Wrong Item Shelf are handled competitively by all three models, reflecting the robustness of large pretrained VLA backbones on simpler manipulation primitives. In contrast, long-horizon sequential tasks—Sorting Packages Continuous (0.16 / 0.03 / 0.00) and Clean the Desktop (0.16 / 0.01 / 0.08)—remain challenging for all models, highlighting the frontier of current policy capabilities and the discriminative value of this suite.
Across all three task suites, $\pi_{0.5}$ consistently achieves the highest average performance, demonstrating advantages in instruction following, perturbation robustness, and manipulation dexterity. The credibility of these cross-model comparisons rests on the reliability of Genie Sim as a benchmark platform, which has been rigorously validated through the sim-to-real experiments in Sec. V-B.
@@ ACKNOWLEDGMENT
We would like to express our sincere gratitude to Linqing Zhong, Sukai Wang and Xiaowei Cai for their assistance with model inference and training; to Haowen Yin, Yanping Zhou, Jiahao Yu, Zengcheng Zhou, and Cheng Ruan for their support with experiment testing; to Chengyue Zhao, Xinke Yu, and Jialu Li for their work on publicity and design; and to Cheng Jing, Haoyu Cao, Chi Zhang, Guangte Xiang, Jun Dai, Jia Zeng for their help with data processing.